### v1

# Towards Explaining Deep Neural Network Compression Through a Probabilistic Latent Space

Abstract
Despite the impressive performance of deep neural networks (DNNs), their computational complexity and storage space consumption have led to the concept of network compression. While DNN compression techniques such as pruning and low-rank decomposition have been extensively studied, there has been insufficient attention paid to their theoretical explanation. In this paper, we propose a novel theoretical framework that leverages a probabilistic latent space of DNN weights and explains the optimal network sparsity by using the information-theoretic divergence measures. We introduce new analogous projected patterns (AP2) and analogous-in-probability projected patterns (AP3) notions for DNNs and prove that there exists a relationship between AP3/AP2 property of layers in the network and its performance. Further, we provide a theoretical analysis that explains the training process of the compressed network. The theoretical results are empirically validated through experiments conducted on standard pre-trained benchmarks, including AlexNet, ResNet50, and VGG16, using CIFAR10 and CIFAR100 datasets. Through our experiments, we highlight the relationship of AP3 and AP2 properties with fine-tuning pruned DNNs and sparsity levels.
1 Introduction
A wide variety of real-world problems are investigated by deep neural networks (DNNs) which enable powerful learning across a wide range of task learning. DNNs have provided remarkable performance in solving complex problems and this has led to significant growth of interest in the area of neural networks. However; the higher the performance, the more the number of parameters, resulting in complexity and space consumption. While DNNs pruning methods [1, 2, 3, 4, 5, 6, 7, 8] compress networks and tackle computational complexity challenges, due to the lack of sufficient theoretical explanation of such methods, it’s difficult to understand their behaviors. In this paper, we propose a novel framework that explains DNN pruning using probabilistic latent spaces of DNN weights and divergence measures. Our main idea is built upon mapping the weight matrix of the layers in both the original network and its corresponding sparse/pruned network to probability spaces from a distribution where the weight matrices are the parameter sets. To clarify our approach: consider a network $F^{(L)}$ with $L$-layers and its sparse version, denoted by $\widetilde{F}^{(L)}$ with parameters $\omega$ and $\widetilde{\omega}=m\odot\omega$, respectively, in which $m$ is a binary mask. Suppose that $f^{(l)}$, filters of the $l$-th layer of $F^{(L)}$ with $M_{l}$ neurons has weight matrix of $\omega^{(l)}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ and $\widetilde{\omega}^{(l)}$ is the weight matrices of $\widetilde{f}^{(l)}$ which is filters of the $l$-th layer of $\widetilde{F}^{(L)}$ again with $M_{l}$ neurons. For clarity, we consistently treat weight matrices as high-dimensional vectors. If $\omega^{(l)}\in\mathbb{R}^{M_{l-1}\times M_{l}}$ represents the weight matrix of layer $l$, its vectorized version is denoted by $\omega^{(l)}\in\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$, where $M_{l}$ is the number of neurons in layer $l$. Note that $\omega^{(l)}$ is referred to as a matrix whenever we use the term "matrix"; otherwise, it is considered as a vector.
In our approach, $\omega^{(l)}$ is projected, using a projection map $\mathcal{P}:\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}\mapsto\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$, to a probability space $\mathcal{Z}$ with random variables $Z$ and with a distribution, denoted by $P_{\omega^{(l)}}$, which has $\omega^{(l)}$ as its parameter set. Similarly, a probability space $\mathcal{\widetilde{Z}}$ with latent variables $\widetilde{Z}$ and distribution $P_{\widetilde{\omega}^{(l)}}$ with $\widetilde{\omega}^{(l)}$ as its parameter set is assigned to $\widetilde{\omega}^{(l)}$ (we illustrate this in Fig. 1). We show that by computing the Kullback-Leibler (KL) divergence between these two distributions, $KL(P_{\omega^{(l)}}||P_{\widetilde{\omega}^{(l)}})$, we explain the behavior of the pruned network and how magnitude-based pruning maintain the task learning performance. We show that as pruned network’s performance convergences to optimal unpruned network performance, the KL divergence between $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$ for given layer $l$ convergences to zero. Note that what distinguishes our method from Bayesian inference is that the networks considered in this paper are non-Bayesian networks, meaning that there is no prior distribution associated with the network parameters. Hence, the distributions are not a distribution of network parameters but rather they are the distribution of defined latent spaces to clarify the network pruning methods.
Why our approach is important? Our new framework enables a deeper understanding of the complex interplay between network pruning and probabilistic distributions, where the weight tensors serve as their key parameters. Our approach effectively explains the sparsity of networks using latent spaces, shedding light on the interpretability of pruned models.
Main Contributions: The key contributions of this study are outlined as follows: (1) We introduce novel projection patterns between layers of DNN and their corresponding compressed counterparts. These patterns allow us to provide an upper bound on the difference between the performance of a network and its pruned version when the projection is applied; (2) We leverage probabilistic latent spaces and employ divergence measures between distributions to explain the sparsity of networks. We show the impact of the differences between two new probabilistic spaces, constructed based on the weights of an original network and its sparse network, on the difference between original and sparse models’ performances; and (3) Through extensive experiments conducted on two well-known CIFAR10 and CIFAR100 datasets, we validate our proposed theoretical framework on multiple benchmarks and showcase how our approach effectively explains the sparsity of networks.
Notations. For clarity and ease of notation, we define all relevant symbols below. As mentioned earlier, we treat $\omega^{(l)}$ as a weight vector by default unless explicitly mentioned as a matrix using the term "matrix". Consider two neural networks, denoted as $F^{(L)}$ and $\widetilde{F}^{(L)}$ with $L$ layers and filters of their $l$-th layer, i.e., $f^{(l)}$ and $\widetilde{f}^{(l)}$. We use $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$ to represent the weight matrices of the $l$-th layer in $F^{(L)}$ and $\widetilde{F}^{(L)}$, respectively. Additionally, $\omega^{*}$ and $\widetilde{\omega}^{*}$ as the optimal weight tensors for $F^{(L)}$ and $\widetilde{F}^{(L)}$. The performance difference between these two networks is quantified by the function $D(F,\widetilde{F})$. To measure the dissimilarity between distributions, we use the Kullback–Leibler divergence denoted as $KL(.||.)$ though the intuition behind our theoretical explanations is based on any distance metrix. Finally, the terms $m$, $\sigma$, and $\mathbb{E}$ correspond to a binary mask, activation function, and expectation operation, respectively.
2 Problem Formulation
Consider input $\mathbf{X}\in\mathcal{X}$ and target $Y\in\mathcal{Y}$, with realization space $\mathcal{X}$ and $\mathcal{Y}$, respectively. Suppose $(\mathbf{X},Y)$ have joint distribution $D$. Training a neural network $F^{(L)}\in\mathcal{F}$ is performed by minimizing a loss function (empirical risk) that decreases with the correlation between the weighted combination of the networks and the label:
$\mathbb{E}_{(\mathbf{X},Y)\sim D}\big{\{}L(F^{(L)}(\mathbf{X}),Y)\big{\}}=-\mathbb{E}_{(\mathbf{X},Y)\sim D}\big{\{}Y\cdot\big{(}b+\sum\limits_{F\in\mathcal{F}}\omega_{F}\cdot F^{(L)}(\mathbf{X})\big{)}\big{\}}.$ (1)
We remove offset $b$ without loss of generality. Define $\ell(\omega):=-\sum\limits_{F\in\mathcal{F}}\omega_{F}\cdot F^{(L)}(\mathbf{X})$, $\omega\in\mathbb{R}^{d}$. Therefore the loss function in (1) becomes $\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{Y\cdot\ell(\omega)\right\}$ with optimal weight solution $\omega^{*}$. Note ${\widetilde{\omega}}^{*}=m^{*}\odot\omega^{*}$ is optimal sparse weight where $m\in[0,1]^{d}$ is the binary mask matrix ($d$ is dimension). The optimal sparse weight is achieved by $(m^{*},\omega^{*}):=\underset{\omega,m}{\arg\min}\,\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{Y\cdot\ell(\omega)\right\}$.
Definition 1.
Analogous Projected Patterns (AP2) - Consider the projected map function $\mathcal{P}~{}:~{}\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}\mapsto\mathbb{R}^{(M_{l-1}\times M_{l})\times 1}$. We say filter $f^{(l)}$ and its sparse version $\widetilde{f}^{(l)}$ have AP2 if for small $\epsilon^{(l)}$, $\left\lVert\mathcal{P}\big{(}f^{(l)}\big{)}-\mathcal{P}\big{(}\widetilde{f}^{(l)}\big{)}\right\rVert^{2}\leq\epsilon^{(l)}$. Equivalently, we say layer $l$ with weight $\omega^{(l)}$ and it’s sparse version with sparse weight $\widetilde{\omega}^{(l)}$ have AP2 if for small $\epsilon^{(l)}$, $\left\lVert\mathcal{P}\big{(}\omega^{(l)}\big{)}-\mathcal{P}\big{(}\widetilde{\omega}^{(l)}\big{)}\right\rVert^{2}\leq\epsilon^{(l)}$ or if $\left\lVert(\omega^{(l)}-\widetilde{\omega}^{(l)})\right\rVert^{2}\leq\epsilon^{(l)}$ when we use identity map as the projection function $\mathcal{P}$.
Figure 1: (Left) Proposed latent space $\mathcal{Z}$ and $\widetilde{\mathcal{Z}}$ from Gaussian distribution with mean parameters $\omega$ and $\widetilde{\omega}$ and covariances $\Sigma$ and $\widetilde{\Sigma}$, respectively: $P_{\omega^{(l)}}=G(\omega^{(l)},\Sigma_{1})$ and $P_{\widetilde{\omega}^{(l)}}=G(\widetilde{\omega}^{(l)},\Sigma_{2})$. (Right) A detailed illustration of maping weights matrices to the latent space as the distribution parameter.
Suppose latent variables $Z\in\mathcal{Z}$ and $\widetilde{Z}\in\widetilde{\mathcal{Z}}$ have distributions $P_{\omega}:=P(.,\omega)$ and $P_{\widetilde{\omega}}:=P(.,\widetilde{\omega})$ with parameter sets $\omega$ and $\widetilde{\omega}$, respectively. $\mathcal{Z}$ and $\widetilde{\mathcal{Z}}$ are latent spaces that could overlap. The Kullback-Leibler (KL) divergence [9] between $P_{\omega}$ and $P_{\widetilde{\omega}}$ is
$KL(P_{\omega}\|P_{\widetilde{\omega}}):=\mathbb{E}_{P_{\omega}}\left[\log\left({P_{\omega}}/{P_{\widetilde{\omega}}}\right)\right].$ (2)
Definition 2.
Analogous-in-Probability Projected Patterns (AP3) - Consider the projection $\mathcal{P}$ that maps $\omega^{(l)}$, the $l$-th layer weight, to the probabilistic space $\mathcal{Z}$ with latent variable ${Z}$ and distribution $P_{\omega^{(l)}}$ with $\omega^{(l)}$ as its parameter set. Similarly projection $\mathcal{P}$ maps sparse layer $l$’s weights, $\widetilde{\omega}^{(l)}$, to space $\widetilde{\mathcal{Z}}$ with latent variable $\widetilde{Z}$ and probability density $P_{\widetilde{\omega}^{(l)}}$. We say layer $l$ and its sparse version have AP3 if for small $\epsilon^{(l)}$, $KL(P_{\omega^{(l)}}\|P_{\widetilde{\omega}^{(l)}})\leq\epsilon^{(l)}$.
Throughout this paper, we assume that all distributions have positive definite covariance matrix.
Note that the projection map in Definitions 1 and 2 are not necessarily the same. To clarify this, we provide an example that shows if layer $l$ and its sparse version have AP2 with an identity projection map, then they have AP3 property with Gaussian projection i.e. projection $\mathcal{P}$ maps $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$ to the latent space when latent variables $Z$ and $\widetilde{Z}$ have Gaussian distributions with means $\omega^{(l)}$ and $\widetilde{\omega}^{(l)}$, respectively. Example - Gaussian Projection: Suppose layer $l$ with weight $\omega^{(l)}$ and its sparse version with weight $\widetilde{\omega}^{(l)}$ have AP3 property with Gaussian projection i.e $P_{\omega^{(l)}}\sim G(\omega^{(l)},\Sigma)$ and $P_{\widetilde{\omega}^{(l)}}\sim G(\widetilde{\omega}^{(l)},\widetilde{\Sigma})$, where $\Sigma$ and $\widetilde{\Sigma}$ are covariance matrices. Using (2) we can write
$KL({P}_{\omega^{(l)}}\|P_{\widetilde{\omega}^{(l)}})=\frac{1}{2}\Big{(}\log\frac{|\Sigma|}{|\widetilde{\Sigma}|}+tr(\Sigma^{-1}\widetilde{\Sigma})+(\omega^{(l)}-\widetilde{\omega}^{(l)})^{T}\Sigma^{-1}(\omega^{(l)}-\widetilde{\omega}^{(l)})-d\Big{)},$ (3)
where $tr$ is the trace of the matrix, $d$ is the dimension and $\log$ is taken to base $e$. Without loss of generality assume that $\Sigma=\widetilde{\Sigma}=\Sigma$, then for constant $C$, (3) is simplified and bounded as $\frac{1}{2}\left(C+(\omega^{(l)}-\widetilde{\omega}^{(l)})^{T}\Sigma^{-1}(\omega^{(l)}-\widetilde{\omega}^{(l)})\right)\leq\frac{1}{2}\left(C+\lambda_{max}\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\right)$. Here $\lambda_{max}$ is the maximum eigenvalue of matrix $\Sigma^{-1}$. Since layer $l$ and its sparse have AP2 property (identity projection), $\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\leq\epsilon^{(l)}$. This implies that layer $l$ and its sparse have AP3 property
$KL({P}_{\omega^{(l)}}\|P_{\widetilde{\omega}^{(l)}})\leq\frac{1}{2}\left(C+\lambda_{max}\epsilon^{(l)}\right),\;\;\hbox{where $C=\log\frac{|\Sigma|}{|\widetilde{\Sigma}|}+tr(\Sigma^{-1}\widetilde{\Sigma})-d$}$ (4)
Note that by considering the assumption of the equality of the covariance matrices, $C=0$, and for smaller $\epsilon^{(l)}$ the bound in (4) becomes tighter.
As observed in the previous example, it was demonstrated that for Gaussian distributions, the AP2 property implies the AP3 property. Building upon this finding, under the same distributional assumptions, based on Eq. 5, we infer that if the latent Gaussian distribution has a covariance matrix $\Sigma$ in which all eigenvalues are greater than or equal to 1, then AP3 property holds true and also implies AP2. Note that $\lambda_{min}$ is the minimum eigenvalues of $\Sigma^{-1}$.
$\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\leq\lambda_{min}\|\omega^{(l)}-\widetilde{\omega}^{(l)}\|^{2}_{2}\leq(\omega^{(l)}-\widetilde{\omega}^{(l)})^{T}\Sigma^{-1}(\omega^{(l)}-\widetilde{\omega}^{(l)}))$ (5)
This means that the AP3 property implies the AP2 property for certain distributions. However, establishing this behavior theoretically for complex multivariate distributions is challenging due to the lack of exact KL values. Consequently, empirical evidence and numerical approximations are often used to study the AP3 property of such distributions.
Definition 3.
[10] (Pinsker’s inequality) If $P$ and $Q$ are two distributions on a measurable space $(X,\mathcal{F})$, then $KL(P\|Q)\geq 2d_{TV}^{2}$.
Definition 4.
[9] Consider a measurable space $(\Omega,\mathcal{A})$ and probability measures $P$ and $Q$ defined on $(\Omega,\mathcal{A})$ on a finite domain $D$. The total variation distance between $P$ and $Q$ is defined as
[…]
Lemma 1.
[11] Let $P$ and $Q$ be two arbitrary distributions in $\mathbb{R}^{d}$ with mean vectors of $m_{P}$ and $m_{Q}$ and covariance matrices of $\Sigma_{P}$ and $\Sigma_{Q}$. Then, by setting constant $a:=m_{P}-m_{Q}$,
$d_{TV}(P,Q)\geq\frac{a^{T}a}{2(tr(\Sigma_{P})+tr(\Sigma_{Q}))+a^{T}a}\;\;\hbox{where $tr(A)$ is the trace of a matrix $A$}.$
Corollary 1.
Let $P$ and $Q$ be two arbitrary distributions in $\mathbb{R}^{d}$ with mean vectors of $m_{P}$ and $m_{Q}$ and covariance matrices of $\Sigma_{P}$ and $\Sigma_{Q}$, respectively. Set constant $a:=m_{p}-m_{Q}$. Using definition 3 and Lemma 1, $KL(P\|Q)\geq 2d^{2}_{TV}(P,Q)\geq 2\left(\frac{a^{T}a}{2(tr(\Sigma_{P})+tr(\Sigma_{Q}))+a^{T}a}\right)^{2}$.
According to the Corollary 1, for any two continuous distributions $P_{\omega^{(l)}}$ and $P_{\widetilde{\omega}^{(l)}}$ in a $d$-dimensional space, both with identity matrix $I\in\mathbb{R}^{d\times d}$ as their covariance matrices, if AP3 is satisfied, then, AP2 holds when $\epsilon<2$. However, it’s important to note that there can be instances where AP3 doesn’t necessarily imply AP2, and this relationship depends on the specific values of $\epsilon$ and the characteristics of the distributions.
Definition 5.
Approximately AP3 in Iteration $t$ Consider the projection $\mathcal{P}$ that maps optimal weight $\omega^{*}$, to the probabilistic space $\mathcal{Z}$ with latent variable ${Z}$ and distribution $P_{\omega^{*}}$ with $\omega^{*}$ as its parameter set . Similarly projection $\mathcal{P}$ maps sparse networks optimal weights, $\widetilde{\omega}^{*}$, to space $\widetilde{\mathcal{Z}}$ with latent variable $\widetilde{Z}$ and distribution $P_{\widetilde{\omega}^{*}}$. Let $\widetilde{\omega}^{t}$ be the weight of pruned network $\widetilde{F}_{t}^{(L)}$ at iteration $t$ in training (with mapped latent space $\widetilde{\mathcal{Z}}$ from distribution $P_{\widetilde{\omega}^{t}}$). We say networks $\widetilde{F}_{t}^{(L)}$ and $F^{(L)}$ have approximately AP3 if for small $\epsilon_{t}$:
$|KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{*}})-KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{t}})|=O(\epsilon_{t}),\;\;\hbox{where as $\epsilon_{t}\rightarrow 0$ when $t\rightarrow\infty$.}$ (6)
3 Theoretical Analysis
In DNNs compression, it is essential to maintain the performance with a core mission of reducing computational complexity and memory usage. However, in many compression methods, after significant computational improvement, we want the performance of the network after compression to be as close as possible to the best performance of the complete (non-compressed) network.
Definition 6.
Performance Difference (PD) - Let $F^{(L)}$ and $\widetilde{F}^{(L)}$ be $L$ layers network and its corresponding compressed network. Using loss function in (1), we define PD by
$D(F^{(L)},\widetilde{F}^{(L)}):=\Big{|}\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{L(F^{(L)}(\mathbf{X}),Y)\right\}-\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{L({\widetilde{F}}^{(L)}(\mathbf{X}),Y)\right\}\Big{|}.$ (7)
We are now in the position to state our first main finding and all proofs in details are provided in Appendix. Prior to introducing the theorems, it’s crucial to underline that each latent space is presupposed to include a covariance matrix among its inherent parameters.
3.1 Performance Difference Bounds
Theorem 1.
Let ${\widetilde{\omega}}^{*}$ be the optimal weight of trained sparse network ${\widetilde{F}}^{(L)}$. Let ${P}_{\omega^{*}}$ be distribution with $\omega^{*}$ as its parameters for trained network $F^{(L)}$ and let ${P}_{{\widetilde{\omega}}^{*}}$ be distribution with ${\widetilde{\omega}}^{*}=m^{*}\odot\omega^{*}$ as its parameters for trained sparse network ${\widetilde{F}}^{(L)}$. If there exists a layer $l\in\{1,\ldots,L\}$ that has AP3 with the corresponding layer in $\widetilde{F}^{(L)}$ i.e. $KL({P}_{\omega^{*(l)}}\|{P}_{\widetilde{\omega}^{*(l)}})\leq\epsilon^{(l)}$, for $20<\epsilon^{(l)}<$ 2, then the performance difference in (7) is bounded by $g(\epsilon^{(l)})$ that is
$D(F^{(L)},\widetilde{F}^{(L)})\leq g(\epsilon^{(l)}),$ (8)
where $g$ is an increasing function of $\epsilon^{(l)}$ i.e. for smaller $\epsilon^{(l)}$, the bound in (8) becomes tighter. Note that, trained sparse network reffers to the pruned network that has undergone fine-tuning after the pruning process.
Theorem 1 is the immediate result of two Lemmas 2 and 3 below together. It is important to note that the function $g(\epsilon^{(l)})$ depends on $\epsilon^{(l)}$ sequentially through $\epsilon^{(L)},\ldots,\epsilon^{(l-1)}$. This means that as in Lemma 3, $\epsilon^{(l+1)}$ is an increasing function of $\epsilon^{(l)}$ for $l,\ldots,L-1$ and according to Lemma 2 the function $g$ is an increasing function of $\epsilon^{(L)}$.
In Lemma 2 and Corollary 2, $\lambda_{\text{max}}$ refers to the maximum eigenvalue of Hessian $\nabla^{2}\ell(\widetilde{\omega}^{*})$. Meanwhile, in Corollary 3, $\lambda_{\text{max}}$ represents the maximum eigenvalue of the inverse covariance matrix $\Sigma^{-1}$. In both corollaries, $\lambda_{\text{min}}$ denotes the minimum eigenvalue of $\Sigma^{-1}$. We also introduce constants $\epsilon^{(L)}$ and $\epsilon^{(l)}$ for reference.
Lemma 2.
Consider a trained network $F^{(L)}$ and its trained sparse version, $\widetilde{F}^{(L)}$ with optimal weights $\omega^{*}$ and $\widetilde{\omega}^{*}$, respectively. Apply the projection map $\mathcal{P}$ on the last layer and let ${P}_{\omega^{*(L)}}$ and ${P}_{{\widetilde{\omega}}^{*(L)}}$ be distributions with $\omega^{*(L)}$ and ${\widetilde{\omega}}^{*(L)}=m^{*(L)}\odot\omega^{*(L)}$ as their parameters for ${F}^{(L)}$ and $\widetilde{F}^{(L)}$, respectively. Note that in $F^{(L)}$ and $\widetilde{F}^{(L)}$, layers $11,\ldots,L-$ 1 are not necessaraly identical. Assume ${\rm Vol}(\mathcal{Z}^{(L)})$ is finite and the last layer of $F^{(L)}$ and $\widetilde{F}^{(L)}$ have AP3 property for $20<\epsilon^{(L)}<$ 2 i.e. $KL({P}_{\omega^{*(L)}}\|{P}_{\widetilde{\omega}^{*(L)}})\leq\epsilon^{(L)}$. Hence, the PD in (8) is bounded by $g^{(L)}(\epsilon^{(L)})$ i.e. $D(F^{(L)},\widetilde{F}^{(L)})\leq g^{(L)}(\epsilon^{(L)})$, where $\widetilde{F}^{(L)}$ is the sparse network with last layer weights $\widetilde{\omega}^{*(L)}$. Here for constant $C$, $g^{(L)}(\epsilon^{(L)}):=\mathbb{E}_{(\mathbf{X},Y)\sim D}\Big{\{}Y\cdot\Big{(}\frac{\lambda_{max}\sqrt{\frac{\epsilon^{(L)}}{2}}C}{2(1-\sqrt{\frac{\epsilon^{(L)}}{2}})}\Big{)}\Big{\}}$, is an increasing function of $\epsilon^{(L)}$ i.e. for smaller $\epsilon^{(L)}$, the bound for $D(F^{(L)},\widetilde{F}^{(L)})$ becomes tighter.
Proof sketch of Lemma 2: While a detailed proof of Lemma 2 is available in Appendix, here’s a concise overview. By considering Equation 1 and by applying Jensen inequality, we have $D(F^{(L)},\widetilde{F}^{(L)})\leq\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{Y\cdot\left|\ell(\omega^{*})-\ell(\widetilde{\omega}^{*})\right|\right\}$. Using both quadratic Taylor approximation for $\ell$ and Corollary 1, $D(F^{(L)},\widetilde{F}^{(L)})\leq\mathbb{E}_{(\mathbf{X},Y)\sim D}\Big{\{}Y\cdot\Big{(}\frac{\lambda_{max}\sqrt{\frac{\epsilon^{(L)}}{2}}C}{2(1-\sqrt{\frac{\epsilon^{(L)}}{2}})}\Big{)}\Big{\}}$, where $C$ is a constant.
Corollary 2.
In Lemma 2, suppose the projection map function is Gaussian projection i.e. $P_{\omega^{*(L)}}\sim G(\omega^{*(L)},\Sigma)$ and $P_{\widetilde{\omega}^{*(L)}}\sim G(\widetilde{\omega}^{*(L)},\Sigma)$ with same positive definite covariance matrices. If the last layer of $F^{(L)}$ and $\widetilde{F}^{(L)}$ have AP3 with bound $\epsilon^{(L)}\geq 0$, then
$D(F^{(L)},\widetilde{F}^{(L)})\leq g_{G}^{(L)}(\epsilon^{(L)}),\;\;\hbox{where$g_{G}^{(L)}(\epsilon^{(L)})=\mathbb{E}_{(\mathbf{X},Y)\sim D}\Big{\{}Y\cdot\Big{(}\frac{\lambda_{max}}{2}\sqrt{\frac{2\epsilon^{(L)}}{\lambda_{min}}}\Big{)}\Big{\}}.$}$ (9)
Lemma 3.
For trained layer $l$ and sparse $l$-layer, let ${P}_{\omega^{*(l)}}$ and ${P}_{{\widetilde{\omega}}^{*(l)}}$ be distributions with $\omega^{*(l)}$ and ${\widetilde{\omega}}^{*(l)}$ as their parameters, respectively. Under the assumptions $KL({P}_{\omega^{*(l)}}\|{P}_{\widetilde{\omega}^{*(l)}})\leq\epsilon^{(l)}$ and ${\rm Vol}(\mathcal{Z})<\infty,$ for $\epsilon^{(l)}\geq 0$, and $\mathbb{E}\big{[}1/({P}_{\widetilde{\omega}^{*(l+1)}})^{2}\big{]}$ being bounded, we have $KL({P}_{\omega^{*(l+1)}}\|{P}_{\widetilde{\omega}^{*(l+1)}})\leq\epsilon^{(l+1)}$, where $\epsilon^{(l+1)}:=~{}C_{l+1}\big{(}\frac{\sqrt{2\epsilon^{(l)}}}{C_{l}{\rm Vol}(\mathcal{Z})}+C_{\sigma^{-1},x}^{(l)}\big{)}$ for constants $C_{l}$, $C_{l+1}$, $C_{\sigma^{-1},x}^{(l)}$ and $\epsilon^{(l+1)}$ is increasing in $\epsilon^{(l)}$.
Corollary 3.
In Lemma 3, suppose the projection map function is Gaussian projection i.e. $P_{\omega^{*(l)}}\sim G(\omega^{*(l)},\Sigma)$ and $P_{\widetilde{\omega}^{*(l)}}\sim G(\widetilde{\omega}^{*(l)},\Sigma)$ with covariance matrix $\Sigma$ for $l=~{}1,\ldots,L-1$. If for $\epsilon^{(l)}\geq 0$, $KL({P}_{\omega^{*(l)}}\|{P}_{\widetilde{\omega}^{*(l)}})\leq\epsilon^{(l)}$ and ${\rm Vol}(\mathcal{Z})<\infty$, then for constants $C_{1}$, $C_{\sigma^{-1},\sigma,x}$, and $K$,
$KL({P}_{\omega^{*(l+1)}}\|{P}_{\widetilde{\omega}^{*(l+1)}})\leq\epsilon^{(l+1)},\;\;\hbox{where \;$\epsilon^{(l+1)}:=\frac{1}{2}\left(C_{1}+\lambda_{max}C^{2}_{\sigma^{-1},\sigma,x}\big{(}\frac{2\epsilon^{(l)}-K}{\lambda_{min}}\big{)}\right)$}.$ (10)
3.2 How AP3 Explains Training Sparse Network?
Theorem 1 links the AP3 property to the performance difference between Networks $F^{(L)}$ and $\widetilde{F}^{(L)}$ on optimal weights $\omega^{*}$ and $\widetilde{\omega}^{*}$. A follow-up question is how the latent space $\mathcal{Z}$ explains the training convergence of compressed network $\widetilde{F}^{(L)}$ to optimal sparsity level with approximately maintaining the performance. To provide an answer to this question we employ Definition 5 and approximately AP3 in iteration $t$ property as stated in the next theorem.
Theorem 2.
Suppose that $P_{\omega^{*}}$ is the distribution of latent variable $Z\in\mathcal{Z}$ for network $F^{(L)}$ and $P_{\widetilde{\omega}^{*}}$ is the distribution of $\widetilde{Z}\in\widetilde{\mathcal{Z}}$ for sparse network $\widetilde{F}^{(L)}$. Let $\widetilde{\omega}^{t}$ be the weight of pruned network $\widetilde{F}_{t}^{(L)}$ after iteration $t$ of training with total $T$ iterations. Under the assumption that $F^{(L)}$ and $\widetilde{F}_{t}^{(L)}$ have approximately AP3 in iteration $t$ i.e. for small $\epsilon_{t}$, $|KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{*}})-KL(P_{\widetilde{\omega}^{*}}\|P_{\widetilde{\omega}^{t}})|=O(\epsilon_{t}),$ then we have $D({F}^{(L)},\widetilde{F}_{t}^{(L)})=O(\epsilon_{t},T)$, equivalently $\left|\mathbb{E}_{(\mathbf{X},Y)\sim D}\left[Y.\ell({\omega}^{*})\right]-\mathbb{E}_{(\mathbf{X},Y)\sim D}\left[Y.\ell(\widetilde{\omega}^{t})\right]\right|=O(\epsilon_{t},T)$.
4 Related Work
Pruning has been a widely-used approach for reducing the size of a neural network by eliminating either unimportant weights or neurons. Existing methods can be categorized into structured and unstructured pruning. structured pruning, studied in [3, 6, 7, 8, 12, 13, 14, 15, 16], involves removing entire structures or components of a network, such as layers, channels, or filters. However, unstructured pruning aims to reduce the number of parameters by zeroing out some weights in the network. Some popular techniques commonly used in unstructured pruning are weight magnitude-based pruning, L1 regularization, Taylor expansion-based pruning, and Reinforcement learning-based pruning  [17, 4, 5, 18, 19, 20, 21].
In this paper, to provide a comprehensive explanation of our findings, we focus on weight magnitude pruning technique, previously examined in [17, 19, 22]. In [17], where weight magnitude pruning is introduced, the authors suggested that a substantial number of network weights can be pruned without causing significant performance deterioration. To establish this notion, they conducted experiments employing models such as AlexNet and VGG16. While their research centers on empirically showcasing the efficacy of such technique, it doesn’t extensively explore the formal theoretical analysis of the subject. Our intent, instead, is to provide a theoretical foundation using probability latent spaces to elucidate the principles of network pruning, particularly weight pruning, and its significant role in preserving accuracy.
5 Experiments
To validate our theoretical analysis empirically, contributing to a more comprehensive understanding of the sparsification techniques in DNNs, we conduct a set of experiments on two common CIFAR10 [23] and CIFAR100 [23] datasets, using three different pre-trained benchmark models: ResNet50 [24] and VGG16 [25], which are pre-trained on ImageNet. The results on AlexNet [26] are provided in Appendix.
5.1 Setup:
Training Phase: We initialized these models’ weights with pre-trained weights as the baseline models. To ensure the results’ significance we performed three trials of training the models. Each trial involved training for 100 epochs using the SGD optimizer with a momentum of 0.9 and learning rates of 0.1 for pre-trained ResNet50 and 0.01 for AlexNet and VGG16. In order to facilitate convergence and enhance training efficiency, we incorporated a learning rate scheduler that dynamically adjusted the learning rate as the model approached convergence.
Pruning Phase: We investigate the effects of weight differences within the last layer of the baseline and its pruned version, using magnitude-based pruning [17], over the PD. We initialize the model with the best parameters obtained from each trial. Three different pruning methods were employed: lowest (removing weights with the lowest magnitudes targeting connections that contributed the least to the model’s performance), highest (removing weights with the highest magnitudes, thus eliminating the most influential connections), and random (a stochastic element by randomly removing weights) magnitude prunings. Different percentages of sparsity level, including 0.1, 0.3, 0.5, and 0.8, are used to provide empirical evidence to support the validity of Lemma 2 for the optimal pruning percentage. We maintained similar experimental setups as before. However, given that we were fine-tuning the trained parameters after pruning, we limited the training to 20 epochs. Note that during each epoch of fine-tuning, the test accuracy of the pruned model is computed for the purpose of comparison with the test accuracy of the original network.
The experimental results section is divided into three main segments, all centered around explaining optimal network sparsity using information-theoretic divergence measures: Section 5.2 investigates the relationship between AP2 and the performance difference of the original network and its sparse version by focusing on the last layer of the network, known as Lemma 2; Section 5.3 extends our analysis to AP3 by considering multivariate T-Student distribution; Finally, section 5.4 involves a crucial comparison between AP2 and AP3, aimed at determining whether AP2 truly constitutes the optimal explanation for PDs or if an opportunity exists to identify parameter sets within distributions that might surpass AP2’s capabilities.
5.2 Multivariate Gaussian Distribution with Diagonal Covariance Matrices
In this section, we aim to validate Lemma 2 by examining the projections of the last layer weights of the networks and their pruned version onto multivariate Gaussian distributions. Specifically, we analyze the mean values of $\omega^{(L)}$ and $\widetilde{\omega}^{(L)}$ in their respective latent spaces, $\mathcal{Z}$ and $\mathcal{\widetilde{Z}}$. We utilize a coefficient of the identity covariance matrix for both latent spaces. By substituting this covariance matrix into the Gaussian distribution example outlined in Section 2, we observe that the KL-divergence value precisely matches AP2 (Gaussian Projection Example). Consequently, the following experiments provide strong evidence for the correctness of Lemma 2 for AP2 across various pre-trained models. Note that in Figs. 2 and 3 Y-axis shows the value of both AP2 and PD computed on testset.
ResNet50: Fig. 2 shows a compelling visual demonstration, reaffirming the correctness of Lemma 2 through the use of AP2 on the pruned ResNet50 model applied to CIFAR10 and CIFAR100 datasets. It is crucial to emphasize that in Fig. 2, while the highest and random magnitude pruning methods consistently validate Lemma 2, the lowest magnitude pruning presents occasional challenges. Specifically, in cases where we prune $30\%$ and $50\%$ of the weights, the performance difference of these two pruning percentages is less than $10\%$. These discrepancies can be attributed to significant differences between theoretical assumptions and experimental observations. For instance, one crucial theoretical assumption in Lemma 2 is the use of optimal weights.
Another factor, especially in lowest magnitude pruning, is the removal of non-essential weights, minimally impacting both AP2 and performance differences. Consequently, this can lead to overlapping results, particularly when stochastic variations come into play. However, despite these occasional overlaps, the close values still provide valuable insights into understanding of optimal network sparsity using information-theoretic divergence measures of the latent distributions.
Further, we conducted another experiment by comparing specific percentages of different pruning methods simultaneously, which is shown in the second row of Fig 2. Despite some variability due to stochastic gradient descent in certain epochs, the expected behavior described in Lemma 2 becomes evident. This signifies that when one pruning method’s KL-divergence decreases relative to another, its PD similarly diminishes. Note that covariance matrices for all methods and different percentages are the same in order to compare them in fairness.
VGG16: To explore the interpretability of network sparsity using the AP2 metric extensively we investigated VGG16 additionally in Fig 3.
We ran additional experiments using Gaussian distribution with non-diagonal covariance matrices, which their KL-divergence serves as AP3. Detailed visual representations are included in Appendix to support our research findings.
5.3 Multivariate T-Student Distribution
Here, we employed the multivariate T-Student distribution, suited for heavy-tailed distributions and outliers in the latent space, enhancing the informativeness and interpretability of AP3 over AP2. We use the Monte Carlo technique to estimate KL-divergence in section 5.2. We set the degrees of freedom in the T-student distribution to $4$, aiming for heavier-tailed distributions and greater dissimilarity from a Gaussian distribution and generated 600,000 data points for improved estimation accuracy. To efficiently estimate KL divergence in high-dimensional spaces, we grouped weight vectors into 100 groups and averaged them, resulting in a 100-dimensional latent variable for KL-divergence estimation. While results for RexNet50 are presented here, experiments on AlexNet and VGG16 are detailed in Appendix. In Fig 4 (Y-axis shows both AP3 values and computed PD values on the test set), we observe that a reduction in KL-divergence of one method compared to another corresponds to a similar pattern in their PDs when meeting the assumptions in Lemma 2. However, some observations deviate from Lemma 2 due to AP3 values exceeding the threshold of $2$.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 2: Comparing AP2 & PD for pruned ResNet50. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
This divergence is evident in the highest magnitude pruning of the first row of Fig. 4 (left) and $80\%$ lowest magnitude in the second row of Fig. 4 (right).
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Comparing AP2 & PD for pruned VGG16. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: Comparing AP3 & PD for pruned ResNet50. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
5.4 Comparing AP2 and AP3
Our goal is to compare AP2 and AP3 metrics to distinguish lowest, highest, and random pruning methods. Both metrics share the same covariance matrix for a fair comparison. We evaluate identical pruning percentages side by side using AP2 and AP3. Consequently, by leveraging AP2, we can gain a more comprehensive insight into which pruning method exhibits superior performance differences. The same comparison for pruned AlexNet is shown in Appendix.
It’s worth mentioning that there is potential for enhancing AP3 over AP2 either by adjusting covariance matrices within the multivariate T-Student distribution framework or by changing the distribution. Careful adjustments could lead to superior performance for AP3 in specific contexts, opening up an intriguing avenue for further investigation.
6 Conclusion
In this study, we investigated network sparsity interpretability by employing probabilistic latent spaces. Our main aim was to explore the concept of network pruning using the Kullback-Leibler (KL) divergence of two probabilistic distributions associated with the original network and its pruned counterpart, utilizing their weight tensors as parameters for the latent spaces. Throughout a series of experiments, we demonstrated the efficacy of two projective patterns, namely AP2 and AP3, in elucidating the sparsity of weight matrices. Our analysis provided valuable explanatory insights into the distributions of pruned and unpruned networks. Remarkably, our results indicated that, despite certain limitations encountered in our experiments compared to theoretical claims, nearly all three pruning methods we employed exhibited alignment with the expectations suggested by Lemma 2 when we use the nearly optimal sparsity level.
Going ahead, we intend to thoroughly investigate the implications of optimal pruning for preserving strong information flow  [27, 28, 29] between layers through the use of probabilistic latent spaces. Our main goal is to substantiate the claim that optimal pruning ensures the retention of highly comparable mutual information between the layers in the pruned and original networks. By this, we intend to use our findings in this work to better tackle the information transfer challenges in domain adaptation problems.
Figure 5: Comparing AP3 and AP2 (Y-axis shows both values) for pruned VGG16 on CIFAR10.
7 Acknowledgement
This work has been supported by NSF CAREER-CCF 5409260; the findings are those of the authors only and do not represent any position of these funding bodies.